OpenStudy (anonymous):

Convert the rectangular coordinates (2,-2) to polar form and find two additional polar representations of this point. Please Help!! Will give MEDAL!!

OpenStudy (anonymous):

Do you know the basics of how to go from rectangular (Cartesian) coordinates to polar?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Alright, instead of having x and y coordinates, polar coordinates uses a radius and an angle. So, out first task is to find the length of the line that connects the origin (0,0) to our point (2,-2)

OpenStudy (anonymous):

Do you know how to find that distance?

OpenStudy (anonymous):

Not sure could you help ???

OpenStudy (anonymous):

Sure. The distance between two points is: \[r = \sqrt{(x_2 - x_1)^2 + (y_2 - y_1)^2}\]

OpenStudy (anonymous):

Your two points are (0,0) and (2,-2). Plug the values into that equation, and find the radius.

OpenStudy (anonymous):

r=2????

OpenStudy (anonymous):

Not quite. \[r = \sqrt{(x_2 - x_1)^2 + (y_2 - y_1)^2} = \sqrt{(2-0)^2 + (-2-0)^2} = \sqrt{(2)^2+(-2)^2} = \sqrt{4+4} = \sqrt{8}\]

OpenStudy (anonymous):

Ack \[=\sqrt{8}\]

OpenStudy (anonymous):

sqrt8=2.828??

OpenStudy (anonymous):

Sounds about right, I'd probably leave it as sqrt(8), but your teacher may want a decimal.

OpenStudy (anonymous):

Our next step is to find the angle. Do you know how to find the angle of a line from the origin to the point (2,-2)?

OpenStudy (anonymous):

Not sure again my teacher didn't explain this

OpenStudy (anonymous):

It's alright :) Ok, so the two components of the point actually tell us the length of the sides of a triangle whose hypotenuse points straight to that point. Our point is (2,-2), which will look like this: